A project currently generates sales of $6.5 million, variable costs equal to 47% of sales, and fixed costs of $1.5 million. The firms tax rate is 35%. What are the effects of the following changes on after-tax profits and cash flows? a. Sales increase from $6.5 million to $8.5 million. What are the effects on after-tax profits and cash flows? (Round your answers to 3 decimal places. Enter your answers in millions of dollars.) After-tax profits increase by Sn/r million After-tax cash flow increase by $689,000 million Variable costs increase to 57% of sales. What are the effects on after-tax profits and cash flows? (Round your answers to 3 decimal places. Enter your answers in millions of dollars.) After-tax profits decrease by Sn/r million. After-tax cash flow decrease by $422,500 million
Answer #1

Sales = $6.5 million

Variable Cost = 47%

Fixed Cost = $1.5 million

Tax Rate = 35%

After tax profit = (1-0.35)((1-0.47)*(6.5) - 1.5)

After Tax Profit = $1.264 millions

After Tax Cashflow = 1.264 + 1.5 = $2.764 million(considering fixed cost as non cash charges)

a.

Sales increased to $8.5 million

After Tax Profit = (1-0.35)((1-0.47)*(8.5) - 1.5)

After tax profit = $1.953 million

So, after tax profit increased by $0.689 million

After -tax cashflow also increased by $0.689 million

b.

Variable cost increased to 57% of sales,

After tax profit = (1-0.35)((1-0.57)*(6.5) - 1.5)

After tax profit = $0.842 million

So after tax profit decreased by $0.422 million

after tax cash flow also decreased by $0.422 million